The width of a rectangle is four feet less than the length. The perimeter is 16 ft. Find the width.
Thx

Answer by rfer(10432)

p=2l+2w
2x-8+2x=16
4x=24
x=6
==========
p=2(6)+2(2)
p=12+4
p=16